# Transfer Learning for Molecular Property Predictions from Small Data Sets

### v1

@@ 3.1 ML benchmarks
PaiNN 0.20 (0.31) 0.56 (0.86)
The remaining models SOAP+NN and SOAP+KRR perform considerably worse on both data sets, in direct comparison SOAP+NN yielding better results for HOPV while SOAP+KRR yields better results for Freesolv. We also note that the MAEs from the different models are much closer for regression on HOPV, where the worst model (SOAP+KRR) has a 63 % higher MAE than the best model (SOAP+SD+GBoost). In contrast, for Freesolv, the worst model (SOAP+NN) has a 375 % higher MAE than the best model (PaiNN), indicating that the learning task is much more complex than for HOPV (cf. also Figure 1. For the following transfer learning with PaiNN, the best remaining model was selected to generate the respective pre-training labels, which is SOAP+SD+GBoost for both data sets.
3.2 Transfer Learning
We use both cheap models from quantum chemistry (LDA-DFT and XTB-DFTB) corrected by simple linear regression, and an ML regressor (SOAP+SD+GBoost, from here on named GBoost) to obtain labels for pre-training the PaiNN models (details in the Methods section). For HOPV, consisting of medium-sized organic molecules, we use the structures of the OE62 data set for pre-training, which contains ca. 62 k diverse structures of medium-sized organic molecules. For Freesolv, which on the other hand consists of small organic molecules, we use the structures of the QM9 data set for pre-training, which contains ca. 132 k organic molecules with up to nine heavy (non-H) atoms. Tables 2 and 3 summarize the results obtained for regression on the HOPV and Freesolv data set, respectively, after PaiNN training from scratch (cf. Table 1), or after pre-training on data labeled with either GBoost, XTB, or LDA, respectively. The learning curves displaying the MAE vs. the number of fine-tuning training points are shown in Figure 2 (here, pre-training is done on the full respective pre-training data set), and the learning curves displaying the MAE vs. the number of pre-training points are shown in Figure 3 (here, fine-tuning is done on the full respective fine-tuning data set), each for the HOPV set on top and the Freesolv set in the bottom.
Table 2: MAEs for HOPV HOMO-LUMO gaps (in eV) from PaiNN obtained after training from scratch, or after pre-training on OE62 data with labels obtained from GBoost, XTB or LDA, respectively. Uncertainties are measured in RMSE (in parentheses). Best MAE results are bold.
@@ 3.2 Transfer Learning
MAE (RMSE) / eV
None 0.20 (0.31)
OE62+GBoost 0.18 (0.25)
OE62+XTB 0.13 (0.17)
OE62+LDA 0.13 (0.19)
Table 3: MAEs for Freesolv solvation energies (in kcal/mol) from PaiNN obtained after training from scratch, or after pre-training on OE62 data with labels obtained from GBoost, XTB or LDA, respectively. Uncertainties are measured in RMSE (in parentheses). Best MAE results are bold.
@@
Figure 2: Fine-tuning learning curves for PaiNN training on HOPV (top) and Freesolv (bottom), after training from scratch (black) or after pre-training on OE62 or QM9 data, respectively, with labels obtained from GBoost (red), XTB (light blue) or LDA-DFT (dark blue). The MAE is plotted against the number of training examples used for fine-tuning (log scale).
@@
Figure 3: Pre-training learning curves for PaiNN training on HOPV (top) and Freesolv (bottom), after training from scratch (black star) or after pre-training on OE62 or QM9 data, respectively, with labels obtained from GBoost (red), XTB (light blue) or LDA-DFT (dark blue). The mean absolute error (MAE) is plotted against the number of training examples used for pre-training.
For HOPV (Table 2), pre-training results in an improved learning performance. The MAE is reduced from 0.20 eV (training from scratch) to 0.18 eV after pre-training on OE62+GBoost, and further to 0.13 eV after pre-training on OE62+XTB or OE62+LDA. The pre-training with GBoost is less effective, probably due to the fact that for this method, some training points were removed from OE62 to avoid creating overly large SOAP descriptors for the ML model. However, the GBoost model might also produce less accurate labels than the QC models corrected by linear regression, since the GBoost MAE (0.19 eV, Table 1) is higher than the MAEs of the LDA or XTB plus linear regression fit on HOPV (MAEs of 0.05 eV for LDA, and 0.12 eV for XTB, respectively).
@@
The fine-tuning learning curves for HOPV (Figure 2, top) show that this improvement of ca. 0.07 eV over training from scratch (black curve) after pre-training on OE62+XTB (light blue) or OE62+LDA (dark blue) is relatively stable, independent of the number of HOPV fine-tuning training points. Only, after pre-training on OE62+GBoost, the improvement obtained from pre-training is much higher when fine-tuning on increasingly many HOPV data points (red curve).
@@
The HOPV pre-training learning curves (Figure 3, top) confirm that for all pre-training data set sizes tested, the labels obtained from the QC models lead to a higher accuracy as compared to the ML-generated labels. Interestingly, for GBoost and LDA labels, increasing the size of the pre-training data set does not monotonically improve the final training results. Instead, there appears to be an optimal training set size at around 20 k pre-training points for GBoost-labeled data and 30 k pre-training points for LDA-labeled data, respectively, after which further improving the pre-training data set size results in higher MAEs after fine-tuning. For XTB-labeled data, a similar optimal pre-training data set size may exist outside the range of the data available here. This indicates that a pre-trained model that is too strongly fitted on the pre-training task may be worse for transfer learning than a more biased model that only learned to grasp high-level concepts from a limited amount of training data.
@@
Regression learning on Freesolv, on the other hand, is not improved by pre-training the PaiNN models (Table 3). Instead, the final training MAEs are increased from 0.56 kcal/mol (training from scratch) to 0.74 kcal/mol (GBoost), 0.60 kcal/mol (XTB), and 0.64 kcal/mol (LDA). The fine-tuning learning curves (Figure 2, bottom) show that an improvement is achieved when fine-tuning is done with very few ($\leq$ ca. 64) data points, QM9+LDA fine-tuning resulting in the largest improvement. However, for larger fine-tuning data set sizes, the results obtained after QM9+XTB or QM9+LDA pre-training are slightly worse as compared to training from scratch, and pre-training with QM9+GBoost results in the worst performance. Furthermore, the pre-training learning curves (Figure 3, bottom) indicate that the loss of performance is approximately independent from the size of the pre-training data set. Moreover, for pre-training data obtained from GBoost and LDA, the final models tend to even decrease in quality after being pre-trained on larger data sets.
@@
As discussed before, the regression task on Freesolv might be more complex as compared to regression on HOPV. As a consequence, the pre-training labels generated by GBoost are less reliable for Freesolv solvation energies than those generated for HOPV HOMO-LUMO gaps (cf. Table 1). Additionally, the methods used for producing the respective pre-training labels (QC methods) are quite similar to the method used for generating HOPV data (standard DFT),30 but quite dissimilar to the method used for generating Freesolv data (”alchemical” calculations based on molecular dynamics simulations, see also the scatter plots in Figure 1).38 These factors may be the reason for why the employed transfer learning strategy does work very well for use with the HOPV data set, but not for Freesolv.
@@
Another hypothesis for why transfer learning might fail in certain scenarios is that the underlying regression task is too easy, and therefore no additional pre-training is needed.21 However, previous results for regression on molecular energies on the QM9 data set (free energy $G$, internal energy $U$, enthalpy $H$) using PaiNN indicated that MAEs as low as ca. 0.14 kcal/mol can be obtained on a sufficiently large data set, which is considerably lower than the MAEs obtained here.6 Also, these quantities ($G$, $U$, $H$) can easily be calculated from only the respective molecular structure, but solvation energy computations additionally need to take into account the shape and intermolecular interactions of the respective solvation shell, thus adding another level of complexity. Therefore, we anticipate that the regression task on Freesolv is not too easy for transfer learning.
@@
In an attempt to obtain pre-trained models that are more robust and therefore allegedly more suitable for transfer learning, we added additional data points to the respective pre-training data sets. For this, we used XTB to generate additional labels for QM9 (HOMO-LUMO gaps) and OE62 (solvation energies). However, we found that pre-training on a joint OE62 and QM9 data set did not improve the final fine-tuning results as compared to pre-training on OE62 or QM9 data only, respectively, for both HOPV and Freesolv training. This is surprising, because NNs usually get more accurate when increasing the amount of training data, and new dissimilar data points can help to create more robust models. Here, instead, the introduction of more data points that are dissimilar to the data contained in the final fine-tuning data set apparently impedes the learning process. In particular, for Freesolv, the underlying problem of having a complex learning task and dissimilar methods for generating pre-training and fine-tuning labels could not be solved by augmenting the pre-training data set. This observation is in line with previous work on transfer learning in chemistry, which stated that the pre-training and fine-tuning data sets should be as closely aligned as possible to obtain the best results.19
@@
Finally, previous works have used discriminative fine-tuning to further optimize transfer learning.29 In this method, the learning rate during fine-tuning is adapted layer-wise, such that the first layers have a much smaller learning rate than the final layers of the model. The idea is that the first layers, which learn to grasp fundamental features of the molecular data, will be sufficiently trained during pre-training, while the last layers, which derive the final predictions, will need more fine-tuning. We implemented discriminative fine-tuning for PaiNN, following the approach outlined by Zaverkin et al.29 Here, the learning rate (lr) for layer n--1 was adjusted as lrn-1 = lrn/5 with lr = 5e-4 being the learning rate of the last layer. However, we found that this method did not yield any improvements over non-discriminative fine-tuning for transfer learning with PaiNN. In contrast, we found that in general, modifications of the learning rate and learning rate decay had a significant impact on the learning performance, and sticking to the parameters identified by the initial hyperparameter sweep continued to yield the best results.
4 Conclusions
@@
This study investigated the use of common ML algorithms for use with small data sets, focusing on HOPV (350 data points) and Freesolv (643 data points). We found that both the MPNN PaiNN, and gradient boosting with regression trees operating on SOAP molecular descriptors concatenated to a set of simple molecular descriptors yield accurate results, clearly outperforming the SOAP+NN and SOAP+KRR architectures.
@@
Furthermore, we presented a method for data-efficient ML on small physical chemistry data sets using transfer learning. The pre-training data was obtained from computationally cheapQC calculations corrected by simple linear regression calibrated on the fine-tuning data set. This method utilizes the fact that many cheap ab initio or semi-empirical techniques yield quantitatively incorrect but qualitatively correct data, which can easily be corrected via linear regression. For the HOPV data set, the introduced method outperforms transfer learning with labels obtained from an ML model. Here, pre-training with labels computed by either XTB or LDA result in very similar learning performances, but XTB computations were much faster than LDA computations. Accordingly, labelling the pre-training data with XTB should be preferred.
@@
However, the same method did not yield improved learning results for regression on the Freesolv data set. This was due to the fact that the underlying learning task is much more complex and the data used for pre-training is too dissimilar to the data contained in Freesolv.
@@
Furthermore, the ML results indicate that pre-training with a limited amount of data points may be beneficial, since the obtained models are slightly more biased and less strongly fitted to the pre-training task. Furthermore, we found that augmenting the pre-training data sets with additional structures that are less similar to the fine-tuning data does not improve the learning performance. Finally, for the models trained here, sticking to the hyperparameters identified by the initial hyperparameter sweeps was more efficient than adapting the learning rates as done in discriminative fine-tuning.
@@
In conclusion, for chemistry ML tasks that use small data sets, an effective transfer learning strategy should use pre-training on molecules that are similar to the molecules of the downstream task, a labelling method that is closely aligned to the method used for labelling the fine-tuning data, and an optimized training set size for the pre-training.
5 Appendix
5.1 Machine Learning Model Hyperparameters
@@
{acknowledgement}
TK and AB acknowledge support from the Helmholtz Einstein International Berlin Research School in Data Science (HEIBRiDS). Computing resources were kindly provided by the Freie Universität Berlin hpc cluster Curta68 and by the Helmholtz-Zentrum Dresden-Rossendorf. The introduction was in part formulated with the help of ChatGPT 3.5.

### v2

@@ III.1 ML benchmarks
PaiNN 0.01 (0.01) 0.56 (0.86)
III.2 Transfer Learning
We use cheap models from quantum chemistry (LDA-DFT and XTB-DFTB) normalized to the same mean (zero) and standard deviation (one) as the fine-tuning data to obtain labels for pre-training the PaiNN models (details in the Methods section). For HOPV, consisting of medium-sized organic molecules, we use the structures of the OE62 data set for pre-training, which contains ca. 62 k diverse structures of medium-sized organic molecules. For Freesolv, which on the other hand consists of small organic molecules, we use the structures of the QM9 data set for pre-training, which contains ca. 132 k organic molecules with up to nine heavy (non-H) atoms. Tables 2 and 3 summarize the results obtained for regression on the HOPV and Freesolv data set, respectively, after PaiNN training from scratch (cf. Table 1), or after pre-training on data labeled with either XTB or LDA, respectively. The learning curves displaying the MAE vs. the number of fine-tuning training points are shown in Figure 2 (here, pre-training is done on the full respective pre-training data set), and the learning curves displaying the MAE vs. the number of pre-training points are shown in Figure 3 (here, fine-tuning is done on the full respective fine-tuning data set), each for the HOPV set on top and the Freesolv set in the bottom. All reported MAEs are obtained as the mean over five independent runs, and the error bars indicate the MAEs’ standard deviations.
Table 2: MAEs for HOPV HOMO-LUMO gaps (in eV) from PaiNN obtained after training from scratch, or after pre-training on OE62 data with labels obtained from XTB and LDA, respectively. Uncertainties are measured in RMSE (in parentheses). Best MAE results are bold.
@@ III.2 Transfer Learning
MAE (RMSE) / meV
None 7.6 (11.9)
OE62+XTB 6.3 (9.7)
OE62+LDA 6.1 (9.6)
Table 3: MAEs for Freesolv solvation energies (in kcal/mol) from PaiNN obtained after training from scratch, or after pre-training on OE62 data with labels obtained from XTB and LDA, respectively. Uncertainties are measured in RMSE (in parentheses). Best MAE results are bold.
@@
Figure 2: Fine-tuning learning curves for PaiNN training on HOPV (top) and Freesolv (bottom), after training from scratch (black) or after pre-training on OE62 or QM9 data, respectively, with labels obtained from XTB (red) and LDA-DFT (blue). The MAE (mean and standard deviation over five runs) is plotted against the number of training examples used for fine-tuning (log scale).
@@
Figure 3: Pre-training learning curves for PaiNN training on HOPV (top) and Freesolv (bottom), after training from scratch (black star) or after pre-training on OE62 or QM9 data, respectively, with labels obtained XTB (red) and LDA-DFT (blue). The MAE (mean and standard deviation over five runs) is plotted against the number of training examples used for pre-training.
@@
For HOPV (Table 2), pre-training results in an improved learning performance. The MAE is reduced from 7.6 meV (training from scratch) to 6.3 meV after pre-training on OE62+XTB, and further to 6.1 meV eV after pre-training on OE62+LDA. The fine-tuning learning curves for HOPV (Figure 2, top) show that this improvement of ca. 1.5 meV over training from scratch (black curve) after pre-training on OE62+XTB (red) or OE62+LDA (blue) is relatively stable, independent of the number of HOPV fine-tuning training points. Also, the standard deviation of the MAEs is reduced after pre-training on either data set.
@@
The HOPV pre-training learning curves (Figure 3, top) confirm that for all pre-training data set sizes tested, the labels obtained from the QC models lead to a higher accuracy as compared to training from scratch. Interestingly, increasing the size of the pre-training data set does not monotonically improve the final training results. Instead, the final training errors after pre-training on 10k data points do not strongly improve when more pre-training data (up to 50k data points) is used. This indicates that a pre-trained model that is too strongly fitted on the pre-training task may be worse for transfer learning than a more biased model that only learned to grasp high-level concepts from a limited amount of training data, which contradicts the findings in the recent study of Hoffmann et al.Hoffmann et al. (2023)
@@
Regression learning on Freesolv, on the other hand, is not improved by pre-training the PaiNN models (Table 3). Instead, the final training MAEs are increased from 0.56 kcal/mol (training from scratch) to 0.61 kcal/mol (XTB), and 0.66 kcal/mol (LDA). The fine-tuning learning curves (Figure 2, bottom) show that no improvement is achieved throughout, independent of the size of the pre-training data set. Also, there is no clear trend in how the size of the pre-training data set influences the result of the subsequent fine-tuning. In addition, the MAE standard deviations increase when pre-training on either data set.
@@
As discussed before, the regression task on Freesolv might be more complex as compared to regression on HOPV. The methods used for producing the respective pre-training labels (QC methods) are quite similar to the method used for generating HOPV data (standard DFT),Lopez et al. (2016) but quite dissimilar to the method used for generating Freesolv data ("alchemical" calculations based on molecular dynamics simulations, see also the scatter plots in Figure 1).Mobley and Guthrie (2014) These factors may be the reason for why the employed transfer learning strategy does work very well for use with the HOPV data set, but not for Freesolv.
@@
Another hypothesis for why transfer learning might fail in certain scenarios is that the underlying regression task is too easy, and therefore no additional pre-training is needed.Sun, Dai, and Yu (2022) However, previous results for regression on molecular energies on the QM9 data set (free energy $G$, internal energy $U$, enthalpy $H$) using PaiNN indicated that MAEs as low as ca. 0.14 kcal/mol can be obtained on a sufficiently large data set, which is considerably lower than the MAEs obtained here.Schütt, Unke, and Gastegger (2021) Also, these quantities ($G$, $U$, $H$) can easily be calculated from only the respective molecular structure, but solvation energy computations additionally need to take into account the shape and intermolecular interactions of the respective solvation shell, thus adding another level of complexity. Therefore, we anticipate that the regression task on Freesolv is not too easy for transfer learning.
@@
In an attempt to obtain pre-trained models that are more robust and therefore allegedly more suitable for transfer learning, we added additional data points to the respective pre-training data sets. For this, we used XTB to generate additional labels for QM9 (HOMO-LUMO gaps) and OE62 (solvation energies). However, we found that pre-training on a joint OE62 and QM9 data set did not improve the final fine-tuning results as compared to pre-training on OE62 or QM9 data only, respectively, for both HOPV and Freesolv training. This is surprising, because NNs usually get more accurate when increasing the amount of training data, and new dissimilar data points can help to create more robust models. Here, instead, the introduction of more data points that are dissimilar to the data contained in the final fine-tuning data set apparently impedes the learning process. In particular, for Freesolv, the underlying problem of having a complex learning task and dissimilar methods for generating pre-training and fine-tuning labels could not be solved by augmenting the pre-training data set. This observation is in line with previous work on transfer learning in chemistry, which stated that the pre-training and fine-tuning data sets should be as closely aligned as possible to obtain the best results.Li et al. (2022a)
@@
Finally, previous works have used discriminative fine-tuning to further optimize transfer learning.Zaverkin et al. (2023) In this method, the learning rate during fine-tuning is adapted layer-wise, such that the first layers have a much smaller learning rate than the final layers of the model. The idea is that the first layers, which learn to grasp fundamental features of the molecular data, will be sufficiently trained during pre-training, while the last layers, which derive the final predictions, will need more fine-tuning. We implemented discriminative fine-tuning for PaiNN, following the approach outlined by Zaverkin et al.Zaverkin et al. (2023) Here, the learning rate (lr) for layer n--1 was adjusted as lrn-1 = lrn/5 with lr = 5e-4 being the learning rate of the last layer. However, we found that this method did not yield any improvements over non-discriminative fine-tuning for transfer learning with PaiNN. In contrast, we found that in general, modifications of the learning rate and learning rate decay had a significant impact on the learning performance, and sticking to the parameters identified by the initial hyperparameter sweep continued to yield the best results.
IV Conclusions
@@
This study investigated the use of common ML algorithms for use with small data sets, focusing on HOPV (350 data points) and Freesolv (643 data points). We found that for these data sets, both the MPNN PaiNN, and gradient boosting with regression trees operating on SOAP molecular descriptors concatenated to a set of simple molecular descriptors yield accurate results, clearly outperforming the SOAP+NN and SOAP+KRR architectures.
@@
Furthermore, we presented a method for data-efficient ML on small physical chemistry data sets using transfer learning which we tested on the two data sets HOPV and Freesolv. The pre-training data was obtained from computationally cheap QC calculations corrected by simple linear regression calibrated on the fine-tuning data set. This method utilizes the fact that many cheap ab initio or semi-empirical techniques yield quantitatively incorrect but qualitatively correct data, which can easily be corrected via linear regression. For the HOPV data set, the introduced method outperforms transfer learning with labels obtained from an ML model. Here, pre-training with labels computed by either XTB or LDA result in very similar learning performances, but XTB computations were much faster than LDA computations. Accordingly, labelling the pre-training data with XTB should be preferred.
@@
However, the same method did not yield improved learning results for regression on the Freesolv data set. We suppose that this may be due to the fact that the underlying learning task is much more complex and the data used for pre-training is too dissimilar to the data contained in Freesolv.
@@
Furthermore, the ML results obtained for HOPV indicate that pre-training with a limited amount of data points may be beneficial, since the obtained models are slightly more biased and less strongly fitted to the pre-training task. Furthermore, we found that augmenting the pre-training data sets with additional structures that are less similar to the fine-tuning data does not improve the learning performance. Finally, for the models trained here, sticking to the hyperparameters identified by the initial hyperparameter sweeps was more efficient than adapting the learning rates as done in discriminative fine-tuning. However, given the limited scope of this study, further research will be needed to investigate whether these findings can be generalized to other data sets as well.
@@
In conclusion, the results obtained from this exploratory study suggest that for chemistry ML tasks that use small data sets, an effective transfer learning strategy should use pre-training on molecules that are similar to the molecules of the downstream task, a labelling method that is closely aligned to the method used for labelling the fine-tuning data, and an optimized training set size for the pre-training. Future studies may investigate whether these findings apply to other data sets as well.
Acknowledgements.
TK and AB acknowledge support from the Helmholtz Einstein International Berlin Research School in Data Science (HEIBRiDS). Computing resources were kindly provided by the Freie Universität Berlin hpc cluster CurtaBennett, Melchers, and Proppe (2020) and by the Helmholtz-Zentrum Dresden-Rossendorf. The introduction was in part formulated with the help of ChatGPT 3.5.
Data Availability Statement
Data openly available in a public repository: The data underlying this study are openly available at https://github.com/ThorrenKirschbaum/TransferLearningPaiNN.
Appendix A Machine Learning Model Hyperparameters